# EconoJax: A Fast & Scalable Economic Simulation in Jax

### v1

@@ Abstract.
Accurate economic simulations often require many experimental runs, particularly when combined with reinforcement learning. Unfortunately, training reinforcement learning agents in multi-agent economic environments can be slow. This paper introduces EconoJax, a fast simulated economy, based on the AI economist. EconoJax, and its training pipeline, are completely written in JAX. This allows EconoJax to scale to large population sizes and perform large experiments, while keeping training times within minutes. Through experiments with populations of 100 agents, we show how real-world economic behavior emerges through training within 15 minutes, in contrast to previous work that required several days. To aid and inspire researchers to build more rich and dynamic economic simulations, we open-source EconoJax on Github at: https://github.com/ponseko/econojax.
Reinforcement Learning, Multi-agent, Economics, JAX
@@ 1. Introduction
Reinforcement learning (Sutton and Barto, 2018) provides an attractive framework for simulating real-world behavior using learned agents in multi-agent environments. Rather than fixing agents to specific strategies, reinforcement learning agents may freely explore the environment and seek economic equilibria without manually implementing certain strategies, which would bias the results. These agent-based models allow for rich, diverse, and interactive simulations in which economic policy questions may be answered (Zheng et al., 2022; Trott et al., 2021).
However, reinforcement learning is known to be sample inefficient (Kaiser et al., 2024), delivering significantly less information per datapoint compared to practices such as supervised or self-supervised learning. Furthermore, in online reinforcement learning, data collection occurs concurrently with the learning process, causing a further increase in computational requirements. Consequently, reinforcement learning agents commonly require days of training, even in simplified, unrealistic simulations.
Over the past decade, we have seen significant performance improvements in deep learning by leveraging efficient GPU’s (Krizhevsky et al., 2012). Although this partly aids deep reinforcement learning agents in their training, real-time environment interactions needed for online learning remain costly. In the last few years, however, tools have been built that allow for faster environment interaction by taking advantage of GPU’s. Of particular interest are WarpDrive (Lan et al., 2021), and JAX (Bradbury et al., 2018), which both provide methods of writing environments to run on GPU accelerated hardware. WarpDrive, however, requires the developer to write specialized CUDA code, which may be difficult to master. In contrast, JAX code is written in regular Python code, albeit with some conditions to the coding style. Importantly, we can also write the reinforcement learning agent code in JAX and unify all code (environment and agent) under one framework. This code can then be compiled by JAX as a single training loop pipeline that operates fully on accelerated hardware, yielding substantial performance improvements, sometimes over a 1000 times  (Lu et al., 2022).
@@
In recent years, the scientific community has made a push to reintroduce existing environments in JAX to address the sample efficiency problem (Matthews et al., 2024; Lange, 2022; Pignatelli et al., 2024; Rutherford et al., 2023; Lu et al., 2022; Bonnet et al., 2024; Freeman et al., 2021; Lu et al., 2023; Frey et al., 2023). However, a similar initiative has not yet emerged for economic agent-based modeling environments. Such an effort could yield significant benefits, as it would allow for richer, more realistic environments, while keeping agent training feasible.
@@
In this work, we introduce EconoJax, a multi-agent economic environment simulating a simplified economy modeled after the AI economist  (Zheng et al., 2022), written in JAX. EconoJax is a vector-based environment with similar functionalities to the AI economist. Population agents gather resources and convert these into coins to maximize their individual utility. The agents interact by trading resources in exchange for coins with other agents. Simultaneously, a government agent sets tax rates at predefined intervals to maximize equality among agents while preserving productivity in the economy.
@@
We experiment with population sizes of 100 agents, which requires roughly 13 minutes of training, and highlight real-world economic behavior emerging in EconoJax, such as realistic tax schedules and the equality-productivity-tradeoff, without manually implementing actions for agents. In comparison, such experiments took many days of training in the original AI economist framework due to a more complex state representation. A rough comparison of training times is shown in Table 1. Due to its increased performance, while still allowing for real-world economic behavior to occur, EconoJax provides a foundation for economic modeling researchers to build richer and more realistic simulations.
@@
The remainder of this paper is organized in the following manner: In the next section, Section 2, we give a brief overview of reinforcement learning, an introduction to JAX, and discuss related work. Section 3 details of EconoJax, including its agents and their respective action and observation spaces. Subsequently, Section 4 displays the outcomes of experiments conducted in EconoJax involving 100 agents. In Section 5 we discuss limitations and future work. Finally, our conclusions are summarized in Section 6.
@@ 2. Preliminaries
Reinforcement Learning is a machine learning method for finding policies that attempt to maximize some reward by interacting with an environment (Sutton and Barto, 2018). Such environments are typically modeled as a Markov Decision Process (MDP) (Bellman, 1957), defined as a tuple ($\mathcal{S}$, $\mathcal{A},p,\mathcal{R},\gamma$). Here, $\mathcal{S}$ is the set of all states, $\mathcal{A}$ the set all possible actions available to the agents, $p$ is the transition function which maps states and actions into a probability distribution of next states $p(s^{\prime}|s,a)$ : $\mathcal{S}x\mathcal{A}x\mathcal{S}\mapsto[0,1]$, $\mathcal{R}$ is the reward function, mapping transitions into rewards $r:\mathcal{S}x\mathcal{A}x\mathcal{S}\mapsto\mathbb{R}$, and $\gamma$ is the discount factor that governs the effect of future rewards on the agent.
@@
JAX is a python library for accelerator oriënted programming (Bradbury et al., 2018), providing convenient function transformations, just-in-time compilation, and a NumPy interface. Just-in-time-compilation support allows JAX functions to be compiled with XLA (Sabne, 2020), such that these functions may efficiently run on platforms such as CPU’s, GPU’s or TPU’s. Furthermore, JAX offers transformations that can easily vectorize, parallelize, or differentiate Python JAX functions. The built-in functionality to differentiate functions (autograd) allows JAX to be used as a deep learning framework, similar to PyTorch and TensorFlow. As such, a large ecosystem of neural network libraries have been built around JAX (Kidger and Garcia, 2021; Hennigan et al., 2020; Godwin* et al., 2020; Heek et al., 2024).
@@
As JAX has increased in popularity, more environments have been reimplemented in the framework. As such, we have JAX based environment suites featuring classic environments such as Gymnax and Jumanji (Lange, 2022; Bonnet et al., 2024), as well as suites that focus on multi-agent reinforcement learning (Rutherford et al., 2023; Willi et al., 2023; de Kock et al., 2023), partial observability (Lu et al., 2023), Physics engines (Freeman et al., 2021), and board games (Koyamada et al., 2023). Furthermore, larger, single environments, rather than suites, are also built in JAX, such as Craftax (Matthews et al., 2024) and Navix (Pignatelli et al., 2024) (MiniGrid).
Agent-based model environments coupled with multi-agent reinforcement learning training could provide interesting insights into (hypothetical) policies and their effect on society. The AI economist (Zheng et al., 2022) demonstrated this by highlighting how reinforcement learning could be used to study the effects of different tax schedules in a simplified economy. They further showed that this would then allow for these tax schedules to be learned as well. This framework was used in a similar fashion to study US state policies for combatting COVID-19 (Trott et al., 2021). More complex economic simulations have recently been built as well, modeling governments, banks, households and firms (Dwarakanath et al., 2024). Other works have studied, using reinforcement learning, how fictitious regions can best negotiate agreements to combat climate change (Zhang et al., 2022).
Recently, an agent-based modeling framework was built in JAX to model the foraging behavior of large populations of organisms (Chaturvedi et al., 2024). However, to the best of our knowledge, no simulator for an economy has yet been released in JAX.
@@
Figure 2. Productivity and equality measured at the end of an episode during training. The shaded areas represent the standard deviation of 15 training runs, in each of which 10 different environments ran in parallel. We see productivity increases over time due to the population agents learning. Introducing taxes manages to significantly improve equality in the population, but trades this for a small bit of productivity.
@@ 3. EconoJax
In this section, we explain the design of EconoJax, an economic simulation modeled after the AI economist (Zheng et al., 2022), created in JAX. The EconoJax source code, along with the experiment code and PPO (Schulman et al., 2017) agent, is available on GitHub: https://github.com/ponseko/econojax, allowing researchers to modify EconoJax for their own research.
@@
EconoJax is a simulation environment for a simple economy modeled after the AI economist framework (Zheng et al., 2022). Similarly to the environments created in the AI economist, EconoJax closely follows the Gymnasium API (Towers et al., 2024) (with modification inspired by Gymnax (Lange, 2022)) and is thus formally an MDP, typically used for reinforcement learning research.
@@
In the default EconoJax implementation, population agents are tasked with maximizing their individual discounted utility. Similar to the AI economist, we model utility as isoelastic utility (Arrow et al., 1974), minus their obtained labor $L$:
(1) $u_{i,t}=\frac{C_{i,t}^{1-\eta}-1}{1-\eta}-L_{i,t}$
@@
Rather than crafting coin, agents are also able to sell their resources to other agents in exchange for coins, allowing for interaction between agents. To facilitate trading, agents can place buy or sell orders on a marketplace. Orders are active for a certain amount of time, and at each timestep all current orders are evaluated. In case there exist buy orders with a value higher or equal to a sell order, these orders are matched – resources and coin are exchanged. In case orders could be matched to multiple other orders, high-value orders (high bids / low asks) are evaluated first and ties are then broken by the age of the order. In order to facilitate trade, coins and resources that are currently used for orders are temporarily placed in agents escrow, rather than their usual inventory.
@@
The government agent sets tax rates in fixed tax brackets with the goal of maximizing equality and productivity. These two measures are seen as important optimization goals for taxation (and the redistribution of those taxes) (Nations Unies, 2013), but this objective may be altered in EconoJax. Tax rates are set every period (default: 100 timesteps). At the end of each period, tax is collected on the earnings from the previous period. The collected taxes are then redistributed uniformly among agents, and new tax rates may be set.
@@
(3) $u_{g}=\text{Eq}(C_{t})\cdot\text{Prod}(C_{t}),$
(4) $\text{Eq}(C_{t})=\omega(1-\text{gini}(C_{t}))+1+\omega,0\leq\omega\leq 1.$
@@
(5) $\text{Prod}(C_{t})=\sum_{i}{C_{i,t}},$
and $\text{gini}(C_{t})$ is the gini index (Gini, 1912), a well-known measure of equality. Similarly to the population agents, the final reward for the government agent is the difference in $u_{g}$ between timesteps, as in Equation 2.
@@
Figure 3. Government returns per episode during training (the combination of Figures 2(b)a,b). The shaded areas represent the standard deviation of 15 training runs, in each of which 10 different environments ran in parallel. Government rewards is measured as productivity $\times$ equality, weighted equally. In the free market setting, productivity is higher, but equality is lower. Near the end of training, the government found appropriate tax rates such that the population has close to optimal productivity, while achieving higher equality.
[…]
Government rewards during training
@@
Table 1. Training times in hours for both EconoJax (averaged over 30 runs) and the AI economist (averaged over 6 runs), with population sizes of 100 and 4 respectively. For the AI economist, we used the available RLlib implementation, using 32 CPU threads. For this comparison we have had to use significantly different hardware, and the two environments cannot be directly compared. In turn, this table only gives a rough indication of the performance benefits of EconoJax. With a simulated population of 25 times larger, EconoJax trains its agents over 750 times faster compared to the AI economist.
@@
pop. size: 4 pop. size: 100
Until convergence [^1] We trained the AI economist agents for 150 million steps. By then, results had not fully converged yet, but training for longer was hindered by large memory requirements. EconoJax was trained for 10 million timesteps. 168.28 0.22
Per million steps 1.12 0.02
@@ 4. Experiments
As shown in Table 1, training EconoJax environments with a population of 100 agents took only about 13 minutes on a single NVIDIA GeForce RTX 3090 GPU. In contrast, reproducing the AI economist results for only 4 agents, using the available RLlib (Liang et al., 2018) implementation, took a full week to complete roughly 150 million training steps on a 16 Intel Xeon E5-2630v3 cores @ 2.40GHz (32 threads), before often running out of memory, disallowing us to complete the full 1 billion training steps required. Due to the difference in hardware and observation space (2D vs 1D observations), this is not a fair direct comparison. However, this $>$750$\times$ performance improvement in a population 25x larger, does provide a clear indication of the substantial decrease in the computational requirements of EconoJax. These reductions allow researchers to experiment with less hardware and enable future expansions of EconoJax for enhanced realism while still keeping training times feasible.
@@
We have experimented in EconoJax with population sizes of 100, far exceeding the population sizes of 4 and 10 experimented with in the AI economist. We trained all our agents using PPO (Schulman et al., 2017), with network sizes of two layers with 128 nodes per layer and tangent activation functions. In these experiments, all population agents shared weights for both the policy and value network. Both population and government agents are able to train simultaneously, in contrast to the two-stage training process required in the AI economist. The results of our experiments are obtained by averaging over 10 different environment seeds during training. Additionally, we repeated training 5 times for 3 different (but similar) population skill distributions and highlight the standard deviation over these 15 runs in shaded areas. The population skill levels are initialized via a Pareto distribution with added random noise. This ensures that most agents are at least reasonably skilled in one action.
@@
Table 2. Hyperparameters for our PPO agents (top), and EconoJax environment parameters used for our experiments (bottom). Both government and population agents used the same hyperparameters during training. Entropy coefficient is linearly annealed over 90% of training steps, whereas learning rate annealling happens over the full training duration. Agent skill levels in our experiments, are set such that most agents are at least reasonably skilled at one action, via a pareto distribution. Random normal noise is then added to the skills levels for a less ”static” population.

### v2

@@ Abstract.
Accurate economic simulations often require many experimental runs, particularly when combined with reinforcement learning. Unfortunately, training reinforcement learning agents in multi-agent economic environments can be slow. This paper introduces EconoJax, a fast simulated economy, based on the AI economist Zheng et al. (2022). EconoJax, and its training pipeline, are completely written in JAX. This allows EconoJax to scale to large population sizes and perform large experiments, while keeping training times within minutes. Through experiments with populations of 100 agents, we show how real-world economic behavior emerges through training within 15 minutes, in contrast to previous work that required several days. We additionally perform experiments in varying sized action spaces to test if some multi-agent methods produce more diverse behavior compared to others. Here, our findings indicate no notable differences in produced behavior with different methods as is sometimes suggested in earlier works. To aid further research, we open-source EconoJax on Github at: https://github.com/ponseko/econojax.
Key words and phrases:
Reinforcement Learning; Multi-agent; Economics; JAX
@@ 1. Introduction
Reinforcement learning Sutton and Barto (2018) provides an attractive framework for simulating real-world behavior using learned agents in multi-agent environments. Rather than fixing agents to specific strategies, reinforcement learning agents may freely explore the environment and seek economic equilibria without manually implementing certain strategies, which would bias the results. These agent-based models allow for rich, diverse, and interactive simulations in which economic policy questions may be answered Zheng et al. (2022); Trott et al. (2021).
However, reinforcement learning is known to be sample inefficient Kaiser et al. (2024), delivering significantly less information per datapoint compared to practices such as supervised or self-supervised learning. Furthermore, in online reinforcement learning, data collection occurs concurrently with the learning process, causing a further increase in computational requirements. Consequently, reinforcement learning agents commonly require days of training, even in simplified, unrealistic simulations.
Over the past decade, we have seen significant performance improvements in deep learning by leveraging efficient GPU’s Krizhevsky et al. (2012). Although this partly aids deep reinforcement learning agents in their training, real-time environment interactions needed for online learning remain costly. In the last few years, however, tools have been built that allow for faster environment interaction by taking advantage of GPU’s. Of particular interest are WarpDrive Lan et al. (2021), and JAX Bradbury et al. (2018), which both provide methods of writing environments to run on GPU accelerated hardware. WarpDrive, however, requires the developer to write specialized CUDA code, which may be difficult to master. In contrast, JAX code is written in regular Python code, albeit with some conditions to the coding style. Importantly, we can also write the reinforcement learning agent code in JAX and unify all code (environment and agent) under one framework. This code can then be compiled by JAX as a single training loop pipeline that operates fully on accelerated hardware, yielding substantial performance improvements, sometimes over a 1000 times  Lu et al. (2022).
@@
In recent years, the scientific community has made a push to reintroduce existing environments in JAX to address the sample efficiency problem Matthews et al. (2024); Lange (2022); Pignatelli et al. (2024); Rutherford et al. (2023); Lu et al. (2022); Bonnet et al. (2024); Freeman et al. (2021); Lu et al. (2023); Frey et al. (2023). However, a similar initiative has not yet emerged for economic agent-based modeling environments. Such an effort could yield significant benefits, as it would allow for richer, more realistic environments, while keeping agent training feasible.
@@
In this work, we introduce EconoJax, a multi-agent economic environment simulating a simplified economy modeled after the AI economist  Zheng et al. (2022), written in JAX. EconoJax is a vector-based environment with similar functionalities to the AI economist. Population agents gather resources and convert these into coins to maximize their individual utility. The agents interact by trading resources in exchange for coins with other agents. Simultaneously, a government agent sets tax rates at predefined intervals to maximize equality among agents while preserving productivity in the economy.
@@
We experiment with population sizes of 100 agents, which requires roughly 13 minutes of training, and highlight real-world economic behavior emerging in EconoJax, such as realistic tax schedules and the equality-productivity-tradeoff, without manually implementing actions for agents. In comparison, such experiments took many days of training in the original AI economist framework due to a more complex state representation. Due to its increased performance, while still allowing for real-world economic behavior to occur, EconoJax provides a foundation for economic modeling researchers to build richer and more realistic simulations.
EconoJax can also flexibly scale to larger population sizes and action spaces. By increasing the number of available resources, agents have the option to specialize in more diverse behavior to optimize their own utility. This makes EconoJax an interesting testbed for multi-agent RL research. We explore the effect of various multi-agent methods on agent behavior, noting that we observe a roughly equal spread in behavior for agents that are trained individually, compared to agents that share all their network parameters.
@@
  • We use EconoJax for additional experiments comparing different multi-agent learning methods in varying sizes of action spaces.
@@
The remainder of this paper is organized in the following manner: In the next section, Section 2, we give a brief overview of reinforcement learning, an introduction to JAX, and discuss related work. Section 3 details of EconoJax, including its agents and their respective action and observation spaces. Subsequently, Section 4 displays the outcomes of experiments conducted in EconoJax involving 100 agents. Next, in Section 5, we experiment with different multi-agent methods to showcase their effect on produced policies in varying action space sizes. In Section 6 we discuss limitations and future work. Finally, our conclusions are summarized in Section 7.
@@ 2. Preliminaries
Reinforcement Learning is a machine learning method for finding policies that attempt to maximize some reward by interacting with an environment Sutton and Barto (2018). Such environments are typically modeled as a Markov Decision Process (MDP) Bellman (1957), defined as a tuple ($\mathcal{S}$, $\mathcal{A},p,\mathcal{R},\gamma$). Here, $\mathcal{S}$ is the set of all states, $\mathcal{A}$ the set all possible actions available to the agents, $p$ is the transition function which maps states and actions into a probability distribution of next states $p(s^{\prime}|s,a)$ : $\mathcal{S}x\mathcal{A}x\mathcal{S}\mapsto[0,1]$, $\mathcal{R}$ is the reward function, mapping transitions into rewards $r:\mathcal{S}x\mathcal{A}x\mathcal{S}\mapsto\mathbb{R}$, and $\gamma$ is the discount factor that governs the effect of future rewards on the agent.
@@
JAX is a python library for accelerator oriënted programming Bradbury et al. (2018), providing convenient function transformations, just-in-time compilation, and a NumPy interface. Just-in-time-compilation support allows JAX functions to be compiled with XLA Sabne (2020), such that these functions may efficiently run on platforms such as CPU’s, GPU’s or TPU’s. Furthermore, JAX offers transformations that can easily vectorize, parallelize, or differentiate Python JAX functions. The built-in functionality to differentiate functions (autograd) allows JAX to be used as a deep learning framework, similar to PyTorch and TensorFlow. As such, a large ecosystem of neural network libraries have been built around JAX Kidger and Garcia (2021); Hennigan et al. (2020); Godwin* et al. (2020); Heek et al. (2024).
@@
As JAX has increased in popularity, more environments have been reimplemented in the framework. As such, we have JAX based environment suites featuring classic environments such as Gymnax and Jumanji Lange (2022); Bonnet et al. (2024), as well as suites that focus on multi-agent reinforcement learning Rutherford et al. (2023); Willi et al. (2023); de Kock et al. (2023), partial observability Lu et al. (2023), Physics engines Freeman et al. (2021), and board games Koyamada et al. (2024). Furthermore, larger, single environments, rather than suites, are also built in JAX, such as Craftax Matthews et al. (2024) and Navix Pignatelli et al. (2024) (MiniGrid).
Agent-based model environments coupled with multi-agent reinforcement learning training could provide interesting insights into (hypothetical) policies and their effect on society. The AI economist Zheng et al. (2022) demonstrated this by highlighting how reinforcement learning could be used to study the effects of different tax schedules in a simplified economy. They further showed that this would then allow for these tax schedules to be learned as well. This framework was used in a similar fashion to study US state policies for combatting COVID-19 Trott et al. (2021). More complex economic simulations have recently been built as well, modeling governments, banks, households and firms Dwarakanath et al. (2024). Other works have studied, using reinforcement learning, how fictitious regions can best negotiate agreements to combat climate change Zhang et al. (2022).
Recently, an agent-based modeling framework was built in JAX to model the foraging behavior of large populations of organisms Chaturvedi et al. (2024). However, to the best of our knowledge, no simulator for an economy has yet been released in JAX.
@@
(c) Government Utility
Figure 2. Productivity and equality measured at the end of an episode during training. The shaded areas represent the standard deviation of 15 training runs, in each of which 10 different environments ran in parallel. We see productivity increases over time due to the population agents learning. Introducing taxes manages to significantly improve equality in the population, but trades this for a small bit of productivity. Government utility (c) is measured as productivity $\cdot$ equality, weighted equally. Near the end of training, the government found appropriate tax rates such that the population has close to optimal productivity, while achieving higher equality.
@@ 3. EconoJax
In this section, we explain the design of EconoJax, an economic simulation modeled after the AI economist Zheng et al. (2022), created in JAX. The EconoJax source code, along with the experiment code and PPO Schulman et al. (2017) agent, is available on GitHub: https://github.com/ponseko/econojax, allowing researchers to modify EconoJax for their own research.
@@
EconoJax is a simulation environment for a simple economy modeled after the AI economist framework Zheng et al. (2022). Similarly to the environments created in the AI economist, EconoJax closely follows the Gymnasium API Towers et al. (2024) (with modification inspired by Gymnax Lange (2022)) and is thus formally an MDP, typically used for reinforcement learning research.
@@
In the default EconoJax implementation, population agents are tasked with maximizing their individual discounted utility. Similar to the AI economist, we model utility as isoelastic utility Arrow (1971), minus their obtained labor $L$:
$u_{i,t}=\frac{C_{i,t}^{1-\eta}-1}{1-\eta}-L_{i,t}$ (1)
@@
Rather than crafting coin, agents are also able to sell their resources to other agents in exchange for coins, allowing for interaction between agents. To facilitate trading, agents can place buy or sell orders on a marketplace. Orders are active for a certain amount of time, and at each timestep all current orders are evaluated. In case there exist buy orders with a value higher or equal to a sell order, these orders are matched – resources and coin are exchanged. In case orders could be matched to multiple other orders, high-value orders (high bids / low asks) are evaluated first. Ties are then first broken by the age of the order and lastly at random. In order to facilitate trade, coins and resources that are currently used for orders are temporarily placed in agents escrow, rather than their usual inventory.
@@
The government agent sets tax rates in fixed tax brackets with the goal of maximizing equality and productivity. These two measures are seen as important optimization goals for taxation (and the redistribution of those taxes) Nations Unies (2013), but this objective may be altered in EconoJax. Tax rates are set every period (default: 100 timesteps). At the end of each period, tax is collected on the earnings from the previous period. The collected taxes are then redistributed uniformly among agents, and new tax rates may be set.
@@
$u_{g}=\text{Eq}(C_{t})\cdot\text{Prod}(C_{t}),$ (3)
$\text{Eq}(C_{t})=\omega(1-\text{gini}(C_{t}))+1+\omega,0\leq\omega\leq 1.$ (4)
@@
$\text{Prod}(C_{t})=\sum_{i}{C_{i,t}},$ (5)
and $\text{gini}(C_{t})$ is the gini index Gini (1912), a well-known measure of equality. Similarly to the population agents, the final reward for the government agent is the difference in $u_{g}$ between timesteps, as in Equation 2.
@@
Table 1. Training times in hours for both EconoJax (averaged over 30 runs) and the AI economist (averaged over 6 runs). For the AI economist, we used the available RLlib implementation, using 32 CPU threads. For this comparison, we have had to use significantly different hardware. Therefor, the two environments cannot be directly compared. In turn, this table only gives a rough indication of the performance benefits of EconoJax. In a same sized population, EconoJax trains its agents over 2.400 times faster compared to the AI economist. This is in large part due to the simplified state representation of EconoJax and the GPU based training powered by JAX.
@@
pop. size: 4 pop. size: 4
Until convergence [^1] We trained the AI economist agents for 150 million steps. By then, results had not fully converged yet, but training for longer was hindered by large memory requirements. EconoJax was trained for 5 million timesteps (10 million in experiments with population sizes of 100). 168.28 0.07
Per million steps 1.12 0.01
@@ 4. Experiments
Training EconoJax environments with a population of 100 agents took only about 13 minutes on a single NVIDIA A100 GPU. In Table 1, we compare EconoJax to the AI Economist by training with only 4 agents. In this setting, EconoJax finishes training in roughly 4 minutes. In contrast, reproducing the AI economist results for the same sized population, using the available RLlib Liang et al. (2018) implementation, took a full week to complete roughly 150 million training steps on a 16 Intel Xeon E5-2630v3 cores @ 2.40GHz (32 threads), before often running out of memory, disallowing us to complete the full 1 billion training steps required. Due to the difference in hardware and observation space (2D vs 1D observations), this is not a fair direct comparison. However, this $>$2400$\times$ performance improvement in a population of the same size, does provide a clear indication of the substantial decrease in the computational requirements of EconoJax. These reductions allow researchers to experiment with less hardware and enable future expansions of EconoJax for enhanced realism while still keeping training times feasible.
@@
We have experimented in EconoJax with population sizes of 100, far exceeding the population sizes of 4 and 10 experimented with in the AI economist. We trained all our agents using PPO Schulman et al. (2017), with network sizes of two layers with 128 nodes per layer and tangent activation functions. In these experiments, all population agents shared weights for both the policy and value network. Both population and government agents are able to train simultaneously, in contrast to the two-stage training process required in the AI economist. The results of our experiments are obtained by averaging over 10 different environment seeds during training. Additionally, we repeated training 5 times for 3 different (but similar) population skill distributions and highlight the standard deviation over these 15 runs in shaded areas. The population skill levels are initialized via a Pareto distribution with added random noise. This ensures that most agents are at least reasonably skilled in one action.
@@
Table 2. Hyperparameters for our PPO agents (top), and EconoJax environment parameters used for our experiments (bottom) in Section 4. Both government and population agents used the same hyperparameters during training. Entropy coefficient is linearly annealed over 90% of training steps, whereas learning rate annealling happens over the full training duration. Agent skill levels in our experiments, are set such that most agents are at least reasonably skilled at one action, via a pareto distribution. Random normal noise is then added to the skills levels for a less ”static” population.